Please help me solve this Quadradic Equation by factoring
3m^=20-7m

Found 2 solutions by richwmiller, blwinbbbles:
Answer by richwmiller(17219)

^ just means that an exponent is coming. you forgot the exponent.
2^3 means 2 cubed or to the third power

Answer by blwinbbbles(106)

I assume that this is 3m^2 = 20-7m so the first thing you must do is get it into standard form of ax^2 + bx + c = 0..to do that you must first add 7m to both sides then subtract 20 from both sides and you will end up with
3m^2+7m-20 = 0 ...now you can factor..use 4 and 5 for 20:
(3m-5)(m+4) = 0 ...now you can say
3m-5=0 then 3m=5 then m=5/3 and/or
m+4=0 then m= -4
go back and use each in the original equation...both work so the answer is both
-4 and 5/3
Hope this helps